Smith’s 17-year, $1,000 par value bonds pay 14 interest annually. The market price of the bonds is $950 and the market's required yield to maturity on a comparable-risk bond is 16 percent. a. What is your yield to maturity on the Smith bonds given the market price of the bonds? b. What should be the value of the Smith bonds given your required rate of return on a comparable-bond? c. You should or should not purchase the Smith bonds because they are currently over-priced or under-priced. PLEASE select one of each.
Answer #1

a) Using financial calculator (BA II plus) to solve for the ytm

Inputs: N= 17

Pv= -950

Pmt= 140 ( 14% × 1,000)

Fv= 1,000

I/y= compute

We get, ytm of the bond as 14.82%

b) Using financial calculator to calculate the price of the bond

Inputs: N= 17

I/y= 16%

Pmt= 140

   Fv= 1,000

Pv= compute

We get, value of the bond as $885.03

c) The Smith's bond is overpriced because of lower yield to maturity than the comparable bond. So purchasing bond from smith would be costly . Therefore, we should not purchase the bond.